# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

### v1

@@ Appendix C Supplementary Results
C.1 Zero-Shot Results of Point Forecasting
@@
Table 7 provides full zero-shot results in Time-Series-Library forecasting benchmark (Wu et al., 2022), including prediction horizons in $\{96,192,336,720\}$. We build Sundial with different model sizes with configurations in Table 5. The context length is set as $2880$. We truncate the model’s predictions for tasks requiring a prediction length less than $F=720$.
@@
We compare most advanced time series foundation models based on their official results, including Time-MoE (Shi et al., 2024b), Timer (Liu et al., 2024a, b), Moirai (Woo et al., 2024), TimesFM (Das et al., 2023b), and Chronos (Ansari et al., 2024). We conduct zero-shot evaluations on datasets that are not included during the pre-training of corresponding models. For each of the evaluated model, we use their maximum input length during inference. The metric (MSE/MAE) is calculated from all predicted windows in the test split of each dataset.
C.2 Zero-Shot Results on GIFT-Eval and FEV Leaderboard
@@
We evaluate our models on GIFT-Eval, a benchmark designed to comprehensively assess forecasting performance across diverse time series. GIFT-Eval includes $23$ datasets covering $144,000$ time series and $177$ million data points, which constitute a total of $97$ forecasting configurations. We use the official evaluation suite established by the research team of SalesForce and report aggregated results in Table 2. We evaluate the probabilistic forecasting performance on the FEV leaderboard, which was originally proposed by Ansari et al. (2024) and established by AutoGluon, which comprises $27$ datasets for zero-shot evaluation. We report aggregated metrics in Figure 5 and assess the inference time in Figure 6. We will release the detailed results by submitting Sundial to their open benchmark in the future.
@@ D.1 Showcases of Sundial
We present zero-shot forecasting showcases on all the datasets from FEV (Ansari et al., 2024) in Figure 10-11, and long-term forecasting datasets (Wu et al., 2022) in Figure 12. By generating multiple predictions with different initial noise, we use the raw prediction to calculate the median and plot the $80\%$ prediction interval.
@@
D.2 Showcases of Pre-trained Generative Forecasters and Deterministic Forecasters
As we introduce generative modeling in time series foundation models, we compare zero-shot forecasting showcases from two types of models in Figure 13-14, including (1) Sundial pre-trained by TimeFlow: as a generative forecaster, it can predict various future possibilities based on the lookback series. (2) Adopting the same backbone and pre-training on TimeBench, a Transformer pre-trained by MSE Loss: as a deterministic forecaster, the model can only output one prediction. During pre-training, the unimodal Gaussian prior specified by MSE can be infeasible to handle large-scale pre-training, manifested as sometimes over-smooth predictions in downstream forecasting tasks. Therefore, we hope this work can inspire future paradigms for pre-training time series foundation models and enhance their applicability to real-world scenarios.
Figure 13: Showcases of Sundial (Left) and a counterpart Transformer pre-trained by MSE Loss (Right).
Figure 14: Showcases of Sundial (Left) and a counterpart Transformer pre-trained by MSE Loss (Right).
@@ Appendix E Limitations
Our models represent an initial effort to incorporate generative modeling into time series foundation models, which enables pre-training on heterogeneous time series without specifying any prior distribution. This approach mitigates mode collapse in representation learning and generates a diverse range of probable predictions compared to previous deterministic forecasters. Despite significant progress, the Sundial family still faces limitations, which tend to generate conservative predictions, such as underestimating trends given a rising slope in context. This hallucination is also observed in Chronos (Ansari et al., 2024). We have yet to determine whether the underlying cause arises from the pre-training distribution or the generative paradigm. This situation may also indicate new opportunities during inference. As we only adopt a naïve sampling strategy that begins with random Gaussian noises, it leaves much room for future improvement in sampling strategy and post-processing.
Another aspect of future development lies in model adaptation. Sundial is pre-trained in a univariate approach to address the discrepancy in variate numbers, which prevents it from explicitly utilizing variate correlations or covariate information. As an increasing number of studies address 2D dimensionality (Liu et al., 2024a; Woo et al., 2024), multivariate pre-training is likely to be conducted for domain-specific time series foundation models. Lastly, while autoregressive models provide flexibility in the input context length, multiple steps of autoregression for long output lengths may still lead to oversmooth predictions and unreliable results. In addition to the forecasting principle emphasizing the importance of complete information (long context), we consider that the output length should be determined based on downstream predictability, where instruction tuning of time series foundation models could serve as a promising solution.
@@ F.1 Real-World Applications
In this work, we present a family of time series foundation models designed to facilitate zero-shot forecasting. Our models employ native tokenization for continuous-valued time series and incorporate a flexible training objective, proposed as TimeFlow Loss, to enable probabilistic forecasting. With an unprecedented distribution learning capability and a trillion-level pre-training scale, our models can be used directly or adapted for various forecasting scenarios, such as energy planning, device maintenance, and financial risk prevention. With multiple predictions that are highly coherent with the input series, our model enhances the reliability of decision-making and streamlines the forecasting pipeline for practitioners. This paper primarily focuses on scientific research and does not present any evident negative social impact.
@@ F.2 Academic Research
We curate TimeBench, a trillion-level time series corpora for pre-training foundation models for time series analysis, which we believe will be beneficial to the research community. Technically, we propose a TimeFlow Loss to facilitate the learning of flexible next-patch distributions. Conditioned on the representations acquired by autoregressive Transformers, our model is endowed with a novel generative capability for probabilistic forecasting. It also enhances the representation learning of Transformers without the need for discrete tokenization. Through pre-training on an unprecedented dataset comprising one trillion time points, we identify subtle scalability bottlenecks that are not solely attributable to architectural design but are predominantly influenced by the training objectives of foundation models. The paradigm of generative modeling applied to autoregressive models may provide valuable insights for the development of continuous-valued foundation models.

### v2

@@ Appendix C Supplementary Results
C.1 Discussion of Mode Collapse
Mode collapse is a failure of representation learning, where a model generates a limited variety of outputs, ignoring the diversity in the training data. For time series foundation models, mode collapse stems from the heterogeneity of the time series distribution, e.g., a similar lookback time series goes into divergent trending. In other words, the semantics of time series patterns are highly unstable. It sometimes leads to over-smooth predictions from models optimized by MSE becuase the results are global-optimal for this loss (See showcases on the right of Figure 14-15). Such a training objective pre-defines a unimodal predictive distribution of data, which struggles to accommodate large-scale datasets like TimeBench.
Our work addresses this phenomenon through generative modeling. Generative forecasters learn flexible distributions without relying on probabilistic priors. We evaluate the distributional metric Continuous Ranked Probability Score (CRPS) to assess the quality of generated predictions across different training objectives. The results indicate that the predictive distribution modeled by TimeFlow is more coherent and diverse compared to alternative training objectives, particularly on the highly diverse GIFT-Eval (Aksu et al., 2024). It validates the effectiveness of TimeFlow in mitigating mode collapse.
Table 7: Zero-shot probabilistic forecasting performance using different training objectives. Averaged CRPS is reported here.
Objective ETTh1 ETTh2 ETTm1 ETTm2 ECL Weather GIFT-Eval
TimeFlow Loss 0.0059 0.0037 0.0057 0.0029 0.0082 0.0021 0.5050
Diffusion Loss 0.0082 0.0053 0.0070 0.0039 0.0095 0.0032 0.5340
MSE Loss 0.0063 0.0040 0.0058 0.0032 0.0080 0.0023 0.6420
C.2 Scaling Behavior Using More Data
We compare Sundial with other time series foundation models that are pre-trained with smaller datasets: Chronos (Ansari et al., 2024) is pre-trained on 94 billion time points, and Moirai is pre-trained on 230 billion time points. As their pre-training datasets are part of the subset of TimeBench, we also conduct pre-training on Sundial using these subsets. As shown in Table 8, these results highlight the scaling behavior of Sundial using larger datasets. Additionally, Sundial still achieves better zero-shot forecasting than its counterpart models with the same pre-training dataset.
Table 8: Zero-shot forecasting performance of models trained on different scales of datasets (measured in time points, pts, and 1B means a billion). We report the averaged results from four prediction lengths $\{96,192,336,720\}$ on Time-Series-Library (Wu et al., 2022).
Model (pts.) Chronos (94B) Moirai (230B) Sundial (94B) Sundial (230B) Sundial (1032B)
Dataset MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.591 0.468 0.417 0.419 0.402 0.429 0.403 0.419 0.411 0.434
ETTh2 0.405 0.410 0.362 0.382 0.377 0.414 0.364 0.398 0.333 0.387
ETTm1 0.645 0.500 0.406 0.385 0.367 0.402 0.352 0.385 0.336 0.377
ETTm2 0.310 0.350 0.311 0.337 0.280 0.341 0.273 0.334 0.258 0.320
ECL 0.214 0.278 0.187 0.274 0.172 0.269 0.171 0.267 0.169 0.265
Weather 0.292 0.315 0.287 0.281 0.254 0.301 0.252 0.297 0.234 0.270
C.3 Performance with Varying Lookback Lengths
Time series foundation models operate independently of training, functioning similarly to statistical methods. Given specific forecasting tasks, one of the most important hyperparameters is the lookback length. Unlike fixed-context models, Sundial offers flexibility for practitioners, allowing the context length to be dynamically adjusted during inference. In Figure 10, we present the performance of Sundial utilizing various lookback lengths. Based on our observations, we contend that performance is largely dependent on the forecasting task itself. Specifically, the size of the lookback window can be tuned to meet the forecasting horizon and data periodicity. Time series foundation models provide a training-free approach for rapid adjustments; still, they should enhance their fundamental long-context capabilities to handle high-frequency data.
Figure 10: Zero-shot forecasting performance using different lookback lengths in $\{480,960,1440,1920,2400,2880\}$. We report the averaged results from four prediction lengths $\{96,192,336,720\}$ on Time-Series-Library (Wu et al., 2022).
C.4 Zero-Shot Results of Point Forecasting
@@
Table 9 provides full zero-shot results on Time-Series-Library forecasting benchmark (Wu et al., 2022), including prediction horizons in $\{96,192,336,720\}$. We build Sundial with different model sizes with configurations in Table 5. The context length is fixed as $2880$. We truncate the model’s predictions for tasks requiring a prediction length less than $F=720$.
@@
We compare the most advanced time series foundation models based on their official checkpoints, including Time-MoE (Shi et al., 2024b), Timer (Liu et al., 2024a, b), Moirai (Woo et al., 2024), TimesFM (Das et al., 2023b), and Chronos (Ansari et al., 2024). We conduct zero-shot evaluations on datasets that are not included during the pre-training of the corresponding models. For each of the evaluated model, we use their maximum input length during inference. Metrics (MSE/MAE) are calculated from all predicted windows in the test split of each dataset following Liu et al. (2024a).
C.5 Zero-Shot Results on GIFT-Eval and FEV Leaderboard
@@
We evaluate our models on GIFT-Eval, a benchmark designed to comprehensively assess forecasting performance across diverse time series. GIFT-Eval includes $23$ datasets covering $144,000$ time series and $177$ million data points, which constitute a total of $97$ forecasting configurations. We use the official evaluation suite established by the research team of Salesforce and report aggregated results in Table 2. We evaluate the performance and inference time on the FEV leaderboard, which was originally proposed by Ansari et al. (2024) and established by AutoGluon, which comprises $27$ datasets for zero-shot evaluation. We report aggregated metrics in Figure 4 and assess the inference time in Figure 5. We released the detailed results by submitting Sundial to their open benchmark [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval..
@@ D.1 Showcases of Sundial
Figure 11-13 present zero-shot forecasting showcases on all the datasets from FEV (Ansari et al., 2024) and TSLib (Wu et al., 2022). By generating $20$ predictions with different initial noise, we estimate the median and $80\%$ prediction interval.
@@
D.2 Showcases of Generative Forecasters and Deterministic Forecasters
As we introduce generative modeling in time series foundation models, we compare zero-shot forecasting showcases from two types of models, including (1) Sundial, a generative forecaster pre-trained by TimeFlow, which can predict multiple future possibilities based on a lookback series. (2) Using the same backbone and TimeBench, a Transformer pre-trained by MSE Loss. As a deterministic forecaster, the model can only output the mean prediction. As depicted in Figure 14-15, the unimodal Gaussian prior specified by MSE can be infeasible to handle large-scale pre-training, manifested as sometimes over-smooth predictions in downstream forecasting tasks. Therefore, we hope this work can inspire future paradigms for pre-training time series foundation models and enhance their applicability to real-world scenarios.
Figure 14: Showcases of Sundial (Left) and the same Transformer backbone pre-trained by MSE Loss (Right). MSE Loss optimizes a deterministic forecaster: given a lookback series, the model can only produce one prediction as the estimation of mean values. This objective may fail to accommodate divergent future variations during large-scale pre-training, leading to mode collapse and over-smooth results (as illustrated in the fourth row). TimeFlow optimizes a generative forecaster: it generates various possibilities observed in the pre-training dataset. Based on these raw predictions, we can estimate the underlying complicated distribution and different statistics. Besides, the greater concentration in generated predictions, the higher the model’s confidence in its predictions.
Figure 15: Supplementary showcases of Sundial (Left) and the same Transformer backbone pre-trained by MSE Loss (Right).
@@ Appendix E Limitations
Our models represent an initial effort to incorporate generative modeling into time series foundation models, which enables pre-training on heterogeneous time series without specifying any prior distribution. This approach mitigates mode collapse in representation learning and generates a diverse range of probable predictions compared to previous deterministic forecasters. Despite significant progress in enlarging model capacity, the Sundial family may still face hallucinations. The performance on very high-frequency data is not guaranteed, since TimeBench contains many middle- and low-frequency time series. Therefore, an important future direction is to generalize Sundial to multi-scale time series. This situation may also indicate new opportunities during inference. As we only adopt a naïve sampling strategy that begins with random Gaussian noise, it leaves much room for future improvement in sampling strategy and post-processing, such as frequency normalization.
Another aspect of future development lies in model adaptation. Sundial is pre-trained in a univariate approach to address the discrepancy in variate numbers, which prevents it from explicitly utilizing variate correlations or covariate information. As an increasing number of studies address 2D dimensionality, multivariate pre-training is likely to be conducted for domain-specific time series foundation models. Lastly, while autoregressive models provide flexibility in the input context length, multiple steps of autoregression may still lead to over-smooth predictions and unreliable results.
@@ F.1 Real-World Applications
In this work, we present Sundial, a family of time series foundation models to facilitate out-of-the-box forecasting. Our models employ native tokenization for continuous-valued time series and incorporate a flexible training objective, proposed as TimeFlow Loss, to enable probabilistic forecasting. With an unprecedented model capacity and a trillion-scale dataset, our models can be used directly or adapted for various forecasting scenarios, such as energy planning, weather forecasting, and financial risk prevention. With multiple predictions generation and a just-in-time inference speed, our model enhances the reliability of decision-making and streamlines the forecasting pipeline for practitioners. This paper primarily focuses on scientific research and does not present any evident negative social impact.
@@ F.2 Academic Research
We curate TimeBench, a trillion-level time series dataset for pre-training foundation models for time series analysis, which we believe will be beneficial to the research community. Technically, we propose a TimeFlow Loss to facilitate the learning of flexible next-patch distributions. Conditioned on the lookback representations acquired by autoregressive Transformers, our model is endowed with a novel generative capability for probabilistic forecasting, enhancing representation learning of Transformers without the need for discrete tokenization. Through pre-training on an unprecedented scale, we identify subtle scalability bottlenecks that are not solely attributable to architectural design but are predominantly influenced by the training objectives of foundation models. The proposed TimeFlow Loss applied to autoregressive and generative models may provide insights for the future development of time series foundation models.